A company pays a dividend of $2 today. Its dividend expects to grow at 5% for the first two years and 6% forever after that and the required rate of return is 9%. What is its stock price today?

STEP BY STEP PLEASE
Answer #1

D1=(2*1.05)=2.1

D2=(2.1*1.05)=2.205

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.205*1.06)/(0.09-0.06)

=77.91

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.1/1.09+2.205/1.09^2+77.91/1.09^2

=$69.36(Approx)